# Oscillations Make Neural Networks Robust to Quantization

### v1

@@ Abstract
Table 3: Cross-bit evaluation of pre-trained ImageNet-1k models fine-tuned on CIFAR-10. Grey background is the target-bit accuracy. Models are trained using different quantization methods (QAT and ours) and bit-widths (ternary, 3-bit, and 4-bit), then evaluated across various bit-widths ranging from ternary to FP32. The grey diagonal shows the results for the bit used during training. Results are means and standard deviations over 5 random seeds. All significant differences between QAT and OsciQuant are shown in bold face.
Section 7 Discussion
@@
We have shown that training with weight oscillations induced via regularization is sufficient in most cases to maintain performance after quantization for ResNet and Tiny ViT. This begs the question whether weight oscillations are also a necessary part of the QAT training process. Indeed, some previous work already points towards this. There are examples claiming that both dampening and/or freezing of oscillations too early in the training process is detrimental to performance after quantization Nagel et al. (2022); Han et al. (2021). And in other case presented in Liu et al. (2023), freezing only the low frequency oscillating weights improves performance. This suggests that weight oscillations are both a necessary and sufficient part of QAT, at least in the early phases of the training process. This further supports our hypothesis that oscillations in QAT have a positive effect on quantization robustness.
Additionally, there might be further benefits to our regularization approach compared to QAT. Our method aims to isolate this crucial part of the training process. This is arguably a more principled approach compared to QAT, where quantization during training combined with STE can lead to a number of side-effects beyond oscillations, which can be highly non-intuitive. We present a simple example in the Appendix Sec. A.1 where replacing a single scalar weight by a product of two scalar weights leads to a non-trivial change in training dynamics when using QAT with the STE.
@@
On the other hand, while it is not clear what the additional effects are during QAT, we do note two consistent deviations from the QAT performance when using our regularization method: QAT outperforms regularization at ternary quantization, whereas our regularization method outperforms QAT in cross-bit accuracy for the ternary and 3-bit case. In A.4, we see how it seems that the cross-bit performance for QAT is upper-bounded by the target-bit performance, which might explain the subpar QAT performance at cross-bit compared to our regularization method which seems bounded by the full precision accuracy. Additionally we can note that while it is stated in Alizadeh et al. (2020); Chmiel et al. (2020) that QAT is not robust to cross-bit quantization, A.4 shows that for some cases the robustness is tied closely to how long the model is trained after the target bit accuracy has converged.
Finally we note in A.2 that in the ResNet-18 model, we see similar results for the hyperparameter sweep for different $\lambda$s, which might suggest that the key for robustness is the presence of oscillations and not their precise nature.
Limitations In our experiments we observed that the robustness to cross-bit quantization improves in later training epochs. In order to further improve robustness one might consider an early stopping criterion that evaluates the performance on cross-bit quantization, which was not done in this work. The same approach could also increase cross-bit quantization robustness of QAT although to a lesser degree than for our method.
We performed our experiments on the CIFAR-10 dataset which might make it more difficult to compare our results with other published works that provide benchmark results for other datasets such as ImageNet-1k.
Section 8 Conclusion
Based on the analysis of a toy model we proposed the hypothesis that weight oscillations during training in deep neural networks make the model robust to quantization.
@@
In Sections 4 and 5 we explain on a toy model how training with QAT and STE leads to oscillations and propose a regularizer that encourages this oscillating behaviour. We confirm that as we increase the strength of the regularization, we empirically observe the appearance of clustering together with oscillations.
@@
Finally we experimentally confirm that the regularizer indeed leads to consistent robustness towards quantization for quantization levels above ternary. Our regularization method achieves comparable performance to QAT above ternary quantization when quantizing to the target-bit seen during optimizing and shows increased robustness compared to QAT in cross-bit quantization with bits greater than the target-bit used in the quantizer during training. All this being evidence of our hypothesis.
@@
Our insights on weight oscillations and their role in quantization robustness open new horizons for model quantization approaches. Our regularization method especially creates interesting possibilities for cross-bit robustness, potentially making our regularization method more appealing than QAT when the goal is to deploy or relase a single set of weights that works across different bit widths or maybe even quantizers. While the regularizer used in our experiments should be viewed as an initial step, we expect that quantization robustness could be further improved by developing oscillation-inducing methods that are adaptive to different learning rates, layer statistics or phases of the training process.
Acknowledgments
@@
Authors thank Tong Chen, Jákup Svøðstein and Sebastian Hammer Eliassen for useful discussions.
@@ Appendix A Appendix
Section A.1 2-layer with single weights
Figure 4: We repeat the toy model experiments, but this time with 2 weights, taking into account that the linear term is no longer 0 in the gradient. We notice at epoch 15 and 18 where the prediction of the quantized model is greater than y, the effect of the terms flip for $w_{2}$.
Consider a linear model $f(x)=w_{2}w_{1}x$, with $w_{1},w_{2}$, input $x$, and target $y\in\mathbb{R}$. The quantized version of this model is defined as $f_{q}(x)=q(w_{2})q(w_{1})x$, where $q(\cdot)$ is the quantizer from Eq. 1. The quadratic loss for the model is given by
@@
$\underbrace{\frac{1}{2}x^{2}\Bigl{(}q(w_{2})^{2}q(w_{1})^{2}-w_{2}^{2}w_{1}^{2}\Bigr{)}}_{\text{quadratic term}}\quad+\quad\underbrace{yx\Bigl{(}w_{2}w_{1}-q(w_{2})q(w_{1})\Bigr{)}}_{\text{linear term}}$
Taking the derivative of $\mathcal{L}$ with respect to $w_{1}$:
@@
We note that the linear term is no longer zero in the gradient and thus for a model consisting of 2 single weight layers we see that there is additional effects from QAT other than oscillations. Additionally because of the non-linearity of the rounding operation, even with the absence of a non-linear activation function, we can no longer reduce the model to a single weight.
Section A.2 Hyperparameters
@@ A.2.1 ResNet-18
Figure 5: Mean over 3 runs of the best validation accuracy for different lambdas. Training a ResNet-18 from scratch. Both ternary and 3-bit is at $10^{-3}$ learning rate and 50% of the data used for training. The plot shows three learning rates, where we for each learning ratue evaluate with the $\lambda$s in the rhs. table. The colored background covers the range between the maximum and minimum value of the quantized validation accuracy with the given $\lambda$s.
In Fig. 5 we see the results of a hyperparameter search over different learning rates and $\lambda$s for a ResNet-18 model. There is a clear trend of seeing the best performance at a learning rate of $10^{-3}$. We note that interestingly there is a comparable performance for a wide range of $\lambda$s, indicating that it is the presence of oscillations which is important for quantization robustness, and not the exact frequency of oscillations.
@@ A.2.2 Tiny ViT
Figure 6: Validation accuracy at different $\lambda$ values and the corresponding best validation accuracies for 3-bit and 2-bit configurations for a single run. Learning rate is set to 1e-4 for fine-tuning. For the 2-bit we test higher $\lambda$ but still see no improvemenet in accuracy. We note how all the $\lambda$s lies close to each other, except for the low of $10^{-2}$
Fig. 6 We note how also the Tiny Vit seems to allow for a wide range of $\lambda$s even though we this time note that $\lambda=1$ performs significantly better than the others.
Section A.3 Epochs and cross-bit robustness
@@
There is an interesting interaction between number of epochs trained and robustness both of our method and QAT. We note how QAT converges first for the target-bit and then over time also converges for the 4 and 8-bit. Additionally we see that QAT seems upper-bounded by the target-bit performance, while this is not the case for our metho. Fig. 7,
Section A.4 Convergence behaviour of Tiny ViT
Figure 8: Regularization with a 3-bit quantizer on a Tiny ViT. We note the peculiar behaviour of the orange line, which is the validation accuracy on the target-bit performance. The performances cycles between $\approx 90\%$ and $10\%$, while the full precision accuracy (The model evaluated without quantized weights) stays some-what stable.
Fig. 8 shows the convergence behaviour of the full precision weights and the quantized weights at target-bit. We note how the Tiny ViT displays a peculiar convergence behaviour, where the accuracy will break, only to go up again. In the Tiny Vit model we quantize the self-attention layers, it is already noted in  Liu et al. [2023] that ViTs are especially vulnerable to quantization of the query and key of a self-attention layer, which might be related to the convergence behaviour we see.
Section A.5 Counting oscillations
Figure 9: The plots shows the weights with a total oscillation count $>0$ in a baseline model, a QAT model and a model regularized with $\lambda=1$ respectively. The y-axis is the percentage of the total weights in the first convolutional layer in a ResNet-18 trained from scratch for 50 epochs. For the baseline model we log the full precision weights at each epoch and then apply the quantizer after training, where as for the regularized model we simply log both the full precision and quantized weights at each epoch.
Using the counting method proposed by Nagel et al. [2022], we notice the biased odd distribution of the bins. In the plot used for the main paper, we count each bin the histogram in iterations of 2. Additionally we note that the proposed method misses the first oscillation count in each weight history.

### v2

@@ 5.2.1 Performance on CIFAR-10
Table 3: Cross-bit evaluation of pre-trained ImageNet-1k models fine-tuned on CIFAR-10. Grey background is the target-bit accuracy. Models are trained using different quantization methods (QAT and ours) and bit-widths (3-bit, and 4-bit), then evaluated across various bit-widths ranging from ternary to FP32. Results are means and standard deviations over 5 random seeds. All significant differences between QAT and Oscillations are shown in bold face.
In this experiment we evaluated the robustness of oscillations-only and QAT towards quantization at bit widths different from the ones used during training.
When using the $\mathcal{R_{\lambda}}$ regularization approach, we applied the regularization term with the training bit width during training and applied PTQ after training at a different quantization level. For QAT we trained using the training bit width and afterwards applied PTQ to the latent weights. For each method we also evaluated the corresponding model without PTQ, directly using the latent weights for inference (reported as FP32).
In Table 3 the results from the experiment are reported. A first observation is that the models produced by $\mathcal{R_{\lambda}}$ regularization consistently achieve nearly full-precision accuracy when quantized at 8-bit or when used without quantization, irrespective of the quantization level used during training. This contrasts with QAT, which produces a viable 8-bit or full-precision model only when trained with at least 4-bit.
Furthermore we see that regularizing using Eq. (19) mostly maintains performance when trained at 3 or 4-bit and quantized at bit level of 3 or 4-bit. QAT also achieves this for Tiny ViT but for ResNet, the accuracy of QAT trained at 3-bit and quantized at other bit widths is barely above random guessing.
Regarding training with ternary quantization, we see that $\mathcal{R_{\lambda}}$ regularization produces models that achieve near full-precision performance for ResNet when quantized at 3-bit or higher. Ternary training for ViT is somewhat peculiar in that it fails to produce a model that is viable when quantized to ternary, whereas the performance of the resulting models starts to show a high level of variability at 4-bit and finally reaches close to full-precision accuracy at 8-bit. In contrast, for both ResNet and ViT, the performance of QAT degrades completely to random guessing when trained with ternary quantization and evaluated at any other quantization level.
5.2.2 Performance on Tiny-ImageNet
Fine-tuning:
Table 4 summarizes the test accuracies for the Tiny-ImageNet dataset. Here we observe the same tendency as with CIFAR-10; oscillations provides a significant increase in accuracy compared to the PTQ baseline. While for the Tiny ViT model $\mathcal{R_{\lambda}}$ regularization is sufficient to recover the quantized accuracy of QAT in both the 3 and 4-bit case, for ResNet18 there is a degradation in accuracy at 3-bit.
Quantization Tiny ViT (FT) ResNet-18 (FT)
Baseline FP32 67.17 $\pm$ 0.67 62.93 $\pm$ 0.55
3-bit PTQ 0.58 $\pm$ 0.16 0.51 $\pm$ 0.03
3-bit QAT 44.29 $\pm$ 0.49 54.08 $\pm$ 0.52
3-bit Oscillations 44.62 $\pm$ 2.47 49.34 $\pm$ 0.76
4-bit PTQ 11.02 $\pm$ 2.11 20.02 $\pm$ 4.80
4-bit QAT 60.61 $\pm$ 0.16 58.31 $\pm$ 0.19
4-bit Oscillations 60.54 $\pm$ 0.37 57.26 $\pm$ 0.33
Table 4: Accuracy on Tiny-ImageNet dataset. Mean and standard deviation is over 3 runs. The models is fine-tuned for 50 epochs on the pretrained ImageNet models. PTQ results is from the FP32 baseline. In both the 3 and 4 bit case, oscillations followed by quantization of the weights matches QAT and is noticeably above the PTQ baseline which has neither oscillations nor QAT.
Robustness to Cross-bit Quantization:
In Table  5 we see the cross-bit results from the Tiny-ImageNet experiments. As with CIFAR-10 we note that the models produced by $\mathcal{R_{\lambda}}$ regularization achieves a better cross-bit performance at bits higher than the target bit. Though we do note a change in the cross-bit behavior; the cross-bit results for 3 and 4-bit is generally lower and not as close tot he FP baseline as in the CIFAR-10 case, yet still there is a significant difference between QAT and $\mathcal{R_{\lambda}}$ regularization.
Model Method $\downarrow$ / Eval. bit $\rightarrow$ FP32 Ternary 3-bit 4-bit 8-bit
[…]
Table 5: Cross-bit evaluation of pre-trained ImageNet-1k models fine-tuned on Tiny-ImageNet. Grey background is the target-bit accuracy. Models are trained using different quantization methods (QAT and ours) and bit-widths (3-bit, and 4-bit), then evaluated across various bit-widths ranging from ternary to FP32. Results are means and standard deviations over 5 random seeds. All significant differences between QAT and Oscillations are shown in bold face.
6 Discussion
@@
We have shown that training with weight oscillations induced via $\mathcal{R_{\lambda}}$ regularization is sufficient in most cases to maintain performance after quantization for ResNet and Tiny ViT. The primary effect of our regularizer is to push weights toward quantization thresholds, which results in weights that are robust to quantization. This “threshold-pushing” dynamic naturally leads to two observable phenomena: weight clustering and oscillations. This begs the question whether weight oscillations are also a necessary part of the QAT training process. Indeed, some previous work already points towards this. There are examples claiming that both dampening and/or freezing of oscillations too early in the training process is detrimental to performance after quantization nagel2022overcoming; ImprovingLowBit. And in other case presented in vitoscillations, freezing only the low frequency oscillating weights improves performance. This suggests that weight oscillations are both a necessary and sufficient part of QAT, at least in the early phases of the training process. This further supports our hypothesis that oscillations in QAT have a positive effect on quantization robustness overall.
@@
Yet we do note deviations from QAT when regularizing with Eq. (19): QAT outperforms $\mathcal{R_{\lambda}}$ regularization at ternary quantization (Appendix A.3), whereas $\mathcal{R_{\lambda}}$ regularization outperforms QAT in cross-bit accuracy for the ternary and 3-bit case. In A.5, we see how it seems that the cross-bit performance for QAT is upper-bounded by the target-bit performance, which might explain the subpar QAT performance at cross-bit compared to $\mathcal{R_{\lambda}}$ regularization which seems bounded by the full-precision accuracy.
Limitations and Future Work
Our theoretical analysis was performed using the toy model in Section 3, and the regularization term is motivated using this analysis. We expect other effects that are not entirely captured by this analysis to be part of QAT. This is explored further in Appendix A.1, where we show how the second term in Eq. (9) is not zero in the gradient when there are multiple layers.
The second term in Eq. (9) is closely related to the oscillations-dampening methods such as the one presented in Equation 6 in nagel2022overcoming, which works by adding a linear pull towards the quantization level. In Appendix A.1 we show how for the linear case with two weights, the second term is no longer zero in the gradient and as such QAT then consists of two components; one that creates oscillations and one that dampens them. From this we can see that oscillations alone do not capture the full dynamics of QAT as analyzed in linear models.
Broader Impact:
The work presented here investigates the oscillations that arise when preparing deep neural networks for quantization using QAT. By characterizing these oscillations, we gain deeper insight into the dynamics of QAT, which can be leveraged to enhance its effectiveness and, consequently, the efficiency of quantized models. Improved efficiency translates into reductions in memory, computation, and energy demands, facilitating the deployment of high-performing quantized networks on low-power edge devices. Some of these applications can have broader positive impact (for example, medical electronics) and also negative impact (for example, surveillance systems or weapons).
7 Conclusions
@@
Based on the analysis of a linear model we hypothesized that weight oscillations during training in deep neural networks make the model robust to quantization akin to QAT. In Sections 3 and 4 we explain how training with QAT and STE leads to oscillations and propose a regularizer that encourages this oscillating behavior. We confirm that as we increase the strength of the regularization, and empirically observe the appearance of clustering together with oscillations.
@@
Finally, we experimentally confirm that the regularizer indeed leads to consistent robustness towards quantization for 3-bit and 4-bit quantization levels. Our oscillations by regularization approach achieves comparable performance to QAT above ternary quantization when quantizing to the target-bit seen during the optimization. Furthermore, we also observe that it shows increased robustness compared to QAT in cross-bit quantization with quantization levels higher than the target-bit used in the quantizer during training. All this is evidence in favor of our hypothesis.
@@
Our insights on weight oscillations and their role in quantization robustness open new horizons for model quantization approaches which usually build on the idea of aligning weights at quantization levels – the opposite of what seems to be the core dynamic in QAT. The regularization approach especially creates interesting possibilities for cross-bit robustness, potentially making the regularization method more appealing than QAT when the goal is to deploy or release a single set of model weights that can work across different bit widths or maybe even quantizers. While the regularizer used in our experiments should be viewed as an initial step, we expect that quantization robustness could be further improved by developing oscillation-inducing methods that are adaptive to different learning rates, layer statistics or phases of the training process.
Acknowledgments:
@@
Authors thank Tong Chen, Jákup Svøðstein, Sophia Wilson, Sebastian Hammer Eliassen, Shivam Adarsh and members of SAINTS Lab for useful discussions throughout.
@@ Appendix A Appendix
A.1 Two-layer with Single Weights
Consider a linear model $f(x)=w_{2}w_{1}x$, with weights $w_{1},w_{2}$, input $x$, and target $y\in\mathbb{R}$. The quantized version of this model is defined as $f_{q}(x)=q(w_{2})q(w_{1})x$, where $q(\cdot)$ is the quantizer from Eq. Equation˜1. The quadratic loss for the model is given by
@@ A.1 Two-layer with Single Weights
$\underbrace{\frac{1}{2}x^{2}\Bigl(q(w_{2})^{2}q(w_{1})^{2}-w_{2}^{2}w_{1}^{2}\Bigr)}_{\text{Quadratic term (Oscillator)}}\quad+\quad\underbrace{yx\Bigl(w_{2}w_{1}-q(w_{2})q(w_{1})\Bigr)}_{\text{Linear term (Oscillation Dampener)}}$ (25)
Taking the derivative of $\mathcal{L}$ w.r.t $w_{1}$:
@@
We note that the linear term is no longer zero in the gradient and thus for a model consisting of two single weight layers we see that there are additional effects from QAT other than oscillations. Furthermore, we would like to clarify that even though the two layered model under consideration does not have explicit non-linear activation functions, it cannot be reduced to a single layer model. This is because of the non-linearity of the rounding operation.
Figure 4: We repeat the toy model experiments, but this time with two weights, taking into account that the linear term is no longer zero in the gradient. We notice at epoch 15 and 18 where the prediction of the quantized model is greater than y, the effect of the terms flip for $w_{2}$.
A.2 Hyperparameters
@@ A.2.1 ResNet-18
[…]
Figure 5: Mean over 3 runs of the best test accuracy for different lambdas. Fine-tuning a pretrained ResNet-18 on CIFAR-10 for 50 epochs. Quantizer is set to 3-bit and $10^{-3}$ learning rate and 100% of the training data is used.
@@
[…]
@@ A.2.2 Tiny ViT
[…]
A.3 Ternary Quantization
Performance comparison for the ternary quantization for different models on CIFAR-10 is reported in Table 6. While both QAT and oscillations improve the PTQ baseline significantly, oscillations degrade compared to QAT, especially for the Tiny ViT. This is in line with previous literature vitoscillations, where transformers have been identified as especially sensitive to oscillations. If oscillations are fundamental to QAT, this still leaves the question of why QAT achieves superior ternary performance. Appendix A.1 offers an indication of the underlying reason. Looking at the gradient of the two-layer toy model, in Eq. 25 we note that the second term is no longer zero in the gradient. This term bears close resemblance to existing formulations of oscillation dampeners, such as in Eq. 6 of nagel2022overcoming. These works by pushing weights towards their nearest quantization level, thereby dampening or nullifying the effect of any oscillator (which pushes weights towards their nearest threshold). We therefore speculate that the main component of QAT is still oscillations, but that QAT also has an inherent dampening mechanism. Given that in a uniform quantizer, the quantization error increases exponentially as we decrease the bits in the quantizer, ternary quantization’s large error magnitude makes the absence of dampening particularly detrimental, resulting in sub-optimal quantized performance compared to QAT.
Quantization MLP5 (FS) ResNet-18 (FS) ResNet-18 (FT) Tiny ViT (FT)
Baseline FP32 51.43 $\pm$ 0.39 83.26 $\pm$ 1.07 83.26 $\pm$ 1.07 96.11 $\pm$ 0.31
Ternary PTQ 10.00 $\pm$ 0.02 10.00 $\pm$ 0.01 10.00 $\pm$ 0.01 9.39 $\pm$ 1.11
Ternary QAT 49.20 $\pm$ 1.34 79.62 $\pm$ 6.42 77.02 $\pm$ 7.57 73.53 $\pm$ 0.77
Ternary Oscillations 36.49 $\pm$ 0.51 61.50 $\pm$ 1.82 44.59 $\pm$ 3.30 13.51 $\pm$ 1.32
Table 6: Performance comparison with ternary quantization on CIFAR-10 dataset. Mean and standard deviation is over 3 runs. The models is fine-tuned for 50 epochs on the pretrained ImageNet models. PTQ results is from the FP32 baseline. For both oscillations only and QAT we see a significant improvement over the PTQ baseline. Yet oscillations degrade significantly compared to QAT, especially for the Tiny Vit. FS: Train from scratch. FT: Fine-tuned.
A.4 Epochs and cross-bit robustness
@@
There is an interesting interaction between number of epochs trained and robustness both of our method and QAT. We note how QAT converges first for the target-bit and then over time also converges for the 4 and 8-bit. Additionally we see that QAT seems upper-bounded by the target-bit performance, while this is not the case for oscillations only as shown in Fig. 8.
A.5 Convergence behavior during oscillations-only optimization
Figure 9: (right) Convergence behavior of ResNet-18. (left) Convergence behavior of a Tiny ViT with regularization with a 3-bit quantizer. Note the peculiar behavior of the orange line, which is the validation accuracy on the target-bit performance. The performances cycles between $\approx 90\%$ and $10\%$, while the full precision accuracy (the model evaluated without quantized weights) stays some-what stable.
Fig. 9 shows the convergence behavior of the full precision weights and the quantized weights at target-bit. We note how the Tiny ViT displays a peculiar convergence behavior, where the accuracy breaks, only to go up again. In the Tiny Vit model we regularize the self-attention layers. It is already noted in  vitoscillations that ViTs are especially vulnerable to oscillations in the query and key of self-attention layers, which might be related to the convergence behavior observed when regularizing with Eq. 19.
A.6 Weight Distribution and Oscillation Frequency
QAT $\lambda=0$ $\lambda=10$
Oscillation Weights Oscillation Weights Oscillation Weights
1_0_conv1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
2_0_conv1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
4_1_conv2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
fc Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Oscillations count and weight distribution side by side across layers and regularization settings. Each pair (left: oscillation, right: weights) corresponds to the same $\lambda$ configuration. Note that weight distribution plots do not share x and y axis.